OpenStudy (anonymous):

In bayes rule, P(th|D)=P(D|th)P(th)/P(D), what are the significant of each of the terms? Thanks

OpenStudy (anonymous):

\[P(\theta|D)=\frac{P(D|\theta)P(\theta)}{P(D)}\]

OpenStudy (anonymous):

\(P(D)\) is the probability of D being true. As an example, D might be obtaining a positive result on a lab test for a disease, and \(P(D)\) would be the probability that a random person taking the test would get a positive result. \(P(\theta)\) is the probability of \(\theta\) being true. As an example, \(\theta\) might be having a disease, and \(P(\theta)\) would be the probability that a random person would be suffering from the disease. \(P(D|\theta)\) is the conditional probability of \(D\) being true given that \(\theta\) is true. In our example this would be the probability that a person suffering from the disease would actually test positive for the disease if they took a lab test. In essence this is a measure of how effective the test is; if the probability \(P(D|\theta)\) is close to 1 then the test is very effective. \(P(\theta|D)\) is the conditional probability of \(\theta\) being true given that \(D\) is true. In our example this would be the probability that a person taking a lab test for a disease and testing positive for it is in fact actually suffering from the disease. (If they're not suffering from the disease then the test has produced a false positive.)

OpenStudy (anonymous):

Hi frankhecker. Thank you for your reply. I should've made it clear that the context actually refers to \[posterior = \frac{likelihood \times prior} {evidence}\]The theta in the original equation actually denotes a random variable rather than an event. And that's where my confusion lies.